A bar (L= 80 cm) moves on two frictional rails where the magnetic field is uniform (B = 0.30 T) and into the paper . If v = 50 cm/s and R = 60 mOmps, determine the magnitude and direction of the magnetic force on the moving bar.
1
Expert's answer
2020-03-30T08:05:27-0400
"F_m=BIL=BL\frac{E_i}{R}=BL\frac{BvL}{R}=\frac{B^2L^2v}{R}"

"F_m=\frac{(0.3)^2(0.8)^2(0.5)}{0.06}=0.48\ N\text{ to the left}"